OpenStudy (anonymous):

if the identity cos^2 x = 1- sin^2 x was verified graphically then the graphs of y= cos^2x and y= sin^2 x would be : a) be horizontally shifted b) identical

OpenStudy (anonymous):

c) be horizontally shifted by pi/2 units relative to each other d)be horizontally shifted by pi units relative to each other e) be vertically shifted by 1 unit relative to eachothe

OpenStudy (anonymous):

@mertsj can u help me out?

OpenStudy (callisto):

I don't know is I can help. Plot the graph y= sin^2 x and y=cos^2x Then you would realise that when y=sin^2 x shifts pi/2 to the left, you get the same graph as y=cos^2 x. That is also true when you compare y=sinx and y=cosx

OpenStudy (anonymous):

so it would be c..

OpenStudy (callisto):

That's what I think.

OpenStudy (anonymous):

thts wht i thought too just wanted to make sure.. thanks. can u see if im right with a couple other questions?

OpenStudy (callisto):

I would try, but I'm not sure if I could help

OpenStudy (anonymous):

10-3sinx=13 + 7sinx the solution to this

OpenStudy (callisto):

10-3sinx=13 + 7sinx -3= 10sinx sin x = -3/10 Now, it's the time to use calculator

OpenStudy (anonymous):

wht abt this root 3 tan x+1 =0

OpenStudy (callisto):

sqrt3 tan x+1 =0 sqrt 3 tanx =-1 tanx = -1/sqrt3 Calculator time again. Remember there are 2 angles